Circular No. 6418

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1996af AND 1996ag IN ANONYMOUS GALAXIES

The Mount Stromlo Abell Cluster Supernova Search Team [L. Germany, Mount Stromlo and Siding Spring Observatories (MSSSO); D. Reiss and C. Stubbs, University of Washington; and B. Schmidt, MSSSO] reports the discovery of two supernovae on CCD images taken by G. Thorpe on the Mount Stromlo Observatory 1.27-m telescope (+ Macho Camera). SN 1996af is located at α = 22^{h}28^{m}31^{s}.13, δ = −68^{o}58′33′′.0 (equinox 2000.0), which is 0′′.7 west and 0′′.3 south of the center of a galaxy situated near Abell 3879; a nearby star has position end figures 44^{s}.02, 01′′.2. SN 1996af is visible on V and R CCD images taken on June 12 and 15 at mag ∼ 19.5 and was not detected to mag 21 on similar frames taken on May 16. SN 1996ag is at α = 21^{h}46^{m}50^{s}.47, δ = −43^{o}41′50′′.4, which is 1′′.7 east and 0′′.4 south of a galaxy situated near Abell 3879; a nearby star has position end figures 45^{s}.09, 43′′.9. SN 1996ag is visible on V and R CCD images taken on June 12 and 15 at mag ∼ 20.5 and was not detected to mag 20.5 on similar frames obtained on May 22.

PKS 0405–385 L. Kedziora-Chudczer, University of Sydney; D. Jauncey, M. Wieringa,

J. Reynolds, and A. Tzioumis, Australia Telescope National Facility; and G. Nicholson, Hartebeesthoek Radio Astronomy Observatory, communicate: “Five days of monitoring of PKS 0405–385 with the Australia Telescope Compact Array between June 7.16 and 11.91 UT revealed rapid, quasiperiodic, quasisinusoidal flux-density variations of large amplitude at 3, 6, 13, and 20 cm. The variations are largest and most rapid at 3 and 6 cm, with changes of up to 1 Jy (from 1.6 to 2.6 Jy at 3 cm) in < 1 hr. At 13 cm, the changes are less rapid (up to 0.4 Jy in 8 hr); at 20 cm, up to 0.2 Jy in 12 hr. Such large-amplitude flux-density variations of an extragalactic radio source is unprecedented on such short timescales. PKS 0405–385 is a quasar of mag 18 at z = 1.285 that is only weakly polarized (3-percent linear polarization at 6 GHz), and it has been identified with an EGRET source. If the observed flux changes are intrinsic to the source, they indicate a brightness temperature above 10^{20} K — some eight orders of magnitude above the inverse-Compton limit for incoherent synchrotron emission.”

COMET C/1996 J1 (EVANS-DRINKWATER) Visual m_{1} estimates by C. S. Morris, Lockwood Valley, CA (0.51-m

reflector): June 8.19 UT, 13.4; 9.19, 13.4; 15.20, 13.4; 16.21, 13.4.

1996 June 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green